# NExON-Bayes: A Bayesian approach to network estimation informed by ordinal covariates

### v1

@@ 3.3 Comparison to existing joint and covariate-dependent approaches
The performance of NExON-Bayes is then compared against two other models that are representative of the methodology incorporated into our proposal. The first is the covariate dependent graphical estimation model (covdepGE) of Helwig et al. (2024) Helwig et al. (2024), which models the precision matrix as a function of the covariate. This approach models the dependence on the covariate via the entries of the precision matrix directly, rather than via the PPI of each edge (as is done in NExON-Bayes). The second approach is the Bayesian Joint Spike-and-Slab Graphical Lasso (SSJGL) of Li et al. (2024) Li et al. (2019), which treats the networks as exchangeable, thereby ignoring any inherent ordering among them. Both of these models are extensions of the graphical spike-and-slab model and use a deterministic inference procedure, as in NExON-Bayes.
@@
We assess performance based on precision, which measures the proportion of predicted edges that are true positives, and recall, which measures the proportion of the true edges that are correctly identified. For NExON-Bayes and covdepGE, we use a $0.5$ threshold on the edge PPI to determine edges, based on the median probability model rule (Barbieri and Berger, 2004). SSJGL does not estimate PPIs and so edges are determined to be present when the corresponding precision matrix entries are non-zero. Figure 2 shows that NExON-Bayes outperforms covdepGE with respect to both precision and recall. The poor performance of covdepGE in this setting can be explained by its overly sparse estimates. By only predicting a low number of edges, the total number of true positive edges will remain low, leading to very poor recall. The precision achieved is also the lowest of any model. Furthermore, covdepGE scales particularly poorly with the number of observations $N$, as it models the precision matrix as a continuous function of the covariate, leading to $|\mathcal{A}|\times N=600$ precision matrix estimates rather than $|\mathcal{A}|=4$ estimates returned by each of the other models.
@@
SSJGL and NExON-Bayes perform much more similarly in comparison, yet SSJGL estimates networks with slightly higher precision but much lower recall in most cases, particularly in networks $(1)$ and $(4)$. This is because SSJGL infers a shared similarity structure across all networks and uses it to inform the network estimates. In this simulated setting, edges that “appear” are mutual to networks $(2),(3)\text{ and }(4)$, whilst edges that “disappear” are mutual to networks $(1),(2)\text{ and }(3)$. Therefore, these edges are often estimated as present in all networks even if they are absent in networks $(1)\text{ and }(4)$ due to the model’s tendency to over-estimate the level of similarity. This is explored further in Section 2.2 of the Supplementary Material. Section 2.3 of the Supplementary Material gives an analysis of computational runtime, which shows that NExON-Bayes achieves the best scalability among the models that estimate multiple networks (see Figure S2 of the Supplementary Material). For up to 6 networks with up to 150 variables ($P\leq 150)$, NExON-Bayes runs in less than 20 minutes. SSJGL is up to an order of magnitude slower than this, whilst covdepGE is excessively slow to the point of infeasibility.
@@
Figure 2: Performance of NExON-Bayes, SSJGL and covdepGE in the simulated scenario where $P=100$, $N=200$ and with $50$ edges “appearing” (absolute value of partial correlation linearly increasing with covariate) and $50$ edges “disappearing” (absolute value of partial correlation linearly decreasing with covariate). Performance is assessed over $100$ replicates for NExON-Bayes and SSJGL and $10$ replicates for covdepGE due to excessive runtime (2020+ hours).
@@ 3.4 Real-World Application
We consider proteomic data from The Cancer Genome Atlas (TCGA) (The Cancer Genome Atlas, 2025), specifically, reverse phase protein array data of patients that have been diagnosed with stage I, II and III breast carcinoma [^2] Stage I BRCA refers to tumours that have not spread beyond the breast and are still small in size, stage II refers to slightly larger tumours that may have spread to 1-3 lymph nodes, whilst stage III cancer refers to cancer that has spread to the “lymph nodes, the skin of the breast or the chest muscle” and is larger in size (Macmillan Cancer Support, 2023). (BRCA). These data have been preprocessed using replicate-based normalisation (Goldman et al., 2020), and contain the expression levels of $P=131$ target proteins for $N_{1}=113$ stage I, $N_{2}=429$ stage II and $N_{3}=140$ stage III individuals. We treat tumour stage as our ordinal covariate capturing patient heterogeneity. Although factors like hormonal status also contribute to heterogeneity, we focus on tumour stage as an ordinal covariate to capture broad trends across disease progression. Our aim is to characterize how tumour severity impacts the proteomic network and whether conditional dependencies between protein pairs strengthen or weaken with tumour progression.
@@
NExON-Bayes’s framework provides unique insights into molecular networks by estimating covariate dependencies. In essence, inspecting the edges corresponding to the largest absolute posterior point estimates of $\beta_{ij}$ could help identify biological processes that characterise tumour progression when used to estimate the proteomic networks of BRCA patients. Figure 3 shows the subnetworks consisting of the edges corresponding to the $50$ smallest negative values and the $50$ largest positive values of the posterior mean estimate of $\beta_{ij}$. To assess whether the borrowing of information facilitated through the inclusion of the edge-specific $\beta_{ij}$ is beneficial in this setting, we then perform a ranked gene set enrichment analysis (GSEA) on the two resulting subnetworks. For each subnetwork, the proteins are ranked by the sum of all adjacent absolute posterior means of $\beta_{ij}$. This way, the proteins are ranked according to the overall strength of their covariate dependencies. To perform a GSEA for the networks estimated by SSL, we need to find an equivalent network that expresses the correlations between precision matrix entry and covariate. To do this, we use ordinary least squares regression to find the regressive coefficients between each precision matrix entry and covariate. The proteins are then ranked similarly as before using sum of adjacent regressive coefficients.
Table 2: Results of the gene set enrichment analysis for the two sub-networks with the largest positive $\beta_{ij}$ estimates and the smallest negative $\beta_{ij}$ estimates using NExON-Bayes and SSL. Gene sets that are uniquely found by NExON-Bayes but not by SSL and vice versa, are listed in the corresponding columns.
$\beta_{ij}>0$ (NExON-Bayes) $\beta_{ij}>0$ (SSL) $\beta_{ij}<0$ (NExON-Bayes) $\beta_{ij}<0$ (SSL)
- Apoptosis modulation and signaling // - Acute Viral Myocarditis - Amytrophic Lateral Sclerosis
@@
- Leptin and adiponectin      AP1 survival signaling
[…]
- ATM signaling
Table 2 shows the results of the GSEA, where many more unique gene sets are identified by NExON-Bayes compared to SSL. Of these, several are of interest with relation to BRCA. For instance, the apoptosis modulation and signaling pathway refers to programmed cell death, a vital process that maintains tissue homeostasis by eliminating damaged or unnecessary cells. In cancer, the evasion of apoptosis allows malignant cells to survive and proliferate uncontrollably. As such, the pathway controlling apoptotic modulation and signaling plays an important role in the progression of cancer (Plati et al., 2011).
[…]
Lastly, the Ataxia Telangiectasia Mutated (ATM) Protein signaling gene set is found uniquely enriched in the positive network of NExON-Bayes. The ATM signaling pathway plays a key role in breast cancer by maintaining genomic stability and influencing treatment responses. As an important regulator of the DNA damage response, ATM mutations can lead to increased susceptibility to malignancy and impact tumor progression (Varadhan et al., 2024).
The ability to uncover these gene sets through encapsulating sample-level heterogeneity effectively in a joint model is an important practical advantage of NExON-Bayes. We have shown that, by taking this joint approach, interesting validated or candidate disease mechanisms are uncovered in a disease which progresses through defined stages. These mechanisms warrant dedicated investigation beyond the scope of this paper.
Finally, to further validate the estimations made by NExON-Bayes, we consult the STRING database (Szklarczyk et al., 2023), which contains protein-protein interactions that have been experimentally validated to a given confidence level. We calculate the percentage of pathways that are found by NExON-Bayes and SSL that have been validated to a 90% confidence threshold and show that NExON-Bayes finds a higher percentage of these validated pathways than SSL for all three BRCA stages (see Section 3 of the Supplementary Material).
@@ 4 Discussion
We have introduced a new method, NExON-Bayes, that jointly estimates multiple graphical models while accounting for sample-level heterogeneity as encapsulated in auxiliary ordinal covariate information. We show that, for such auxillary data, NExON-Bayes’ formulation better captures sample-level heterogeneity than network estimation models that use continuous covariates (covdepGE) or treat pre-defined groups of samples in an exchangeable manner, without resorting to covariate information (SSJGL). Moreover, thanks to its efficient deterministic inference procedure, our approach is more scalable and better suited to the analysis of current real data settings. Applying our novel methodology to the TCGA breast cancer data set reveals new molecular pathways that characterize tumour stage progression, which could lead to new hypotheses about plausible oncogenic mechanisms.
Several avenues for future work could be considered. First, the current prior formulation assumes a linear relationship between the ordinal variable and the probability of edge inclusion – relaxing this linearity assumption through an alternative monotonic parametric formulation may enhance both the model performance and its biological relevance. Second, as is typical for joint network estimation models, unequal sample sizes can lead to varying levels of sparsity in the estimated networks, reflecting differences in statistical power. This behaviour can in fact be desirable, as smaller sample sizes naturally yield sparser estimates, reducing the risk of false positives. However, if the goal is to isolate true biological variation independent of sample size effects, this pattern can be adjusted for — though doing so will require methodological advances beyond current approaches (Das et al., 2020).
@@
This research was supported by the Lopez–Loreta Foundation (J.F., H.R., C.L., X.X.). C.L. has also received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No.101126636. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.

### v2

@@ 3.3 Comparison to existing joint and covariate-dependent approaches
The performance of NExON-Bayes is then compared against three other models that are representative of, or similar to, the methodology incorporated into our proposal. The first is the covariate dependent graphical estimation model (covdepGE) of helwig2024algorithm, which models the precision matrix as a function of the covariate. This approach models the dependence on the covariate via the entries of the precision matrix directly, rather than via the PPI of each edge (as is done in NExON-Bayes). The second approach is the Bayesian Joint Spike-and-Slab Graphical Lasso (SSJGL) of Li2019, which treats the networks as exchangeable, thereby ignoring any inherent ordering among them. Both of these models are extensions of the graphical spike-and-slab model and use a deterministic inference procedure, as in NExON-Bayes. Lastly, the third is the multiple graphical horseshoe (mGHS) of Busatto02102025, which treats networks exchangeably and uses a horseshoe prior on precision matrix entries along with an MCMC algorithm.
@@
We principally assess performance based on precision, which measures the proportion of predicted edges that are true positives, and recall, which measures the proportion of the true edges that are correctly identified. For NExON-Bayes and covdepGE, we use a $0.5$ threshold on the edge PPI to determine edges, based on the median probability model rule (barbieri2004optimal). SSJGL does not estimate PPIs and so edges are determined to be present when the corresponding precision matrix entries are non-zero. For mGHS, we follow the $t^{\alpha}$ thresholding rule introduced by the authors (Busatto02102025), which uses a Metropolis-within-Gibbs step to update the threshold parameters. Figure 3 shows the precision and recall of the estimates of each model for each network; for full results that include MCC, accuracy and sparsity, see Supplementary Material 2.3 Table S1. NExON-Bayes outperforms covdepGE with respect to precision, recall, MCC and accuracy. The poor performance of covdepGE in this setting can be explained by its overly sparse estimates. By only predicting a low number of edges, the total number of true positive edges will remain low, leading to very poor recall. The precision achieved is also the lowest of any model. Furthermore, covdepGE scales particularly poorly with the number of observations $N$, as it models the precision matrix as a continuous function of the covariate, leading to $|\mathcal{A}|\times N=600$ precision matrix estimates rather than $|\mathcal{A}|=4$ estimates returned by each of the other models.
@@
SSJGL and NExON-Bayes perform much more similarly in comparison, yet SSJGL estimates networks with slightly higher precision but much lower recall in most cases, particularly in networks $(1)$ and $(4)$. This is because SSJGL infers a shared similarity structure across all networks and uses it to inform the network estimates. In this simulated setting, edges that ‘appear’ are mutual to networks $(2),(3)\text{ and }(4)$, whilst edges that ‘disappear’ are mutual to networks $(1),(2)\text{ and }(3)$. Therefore, these edges are often estimated as present in all networks even if they are absent in networks $(1)\text{ and }(4)$ due to the model’s tendency to over-estimate the level of similarity. This is explored further in Supplementary Material 2.2.
mGHS exhibits the opposite problem to covdepGE, in that it estimates very dense networks, indicating poor convergence. It achieves the highest recall of any model, which is a misleading result common for such overly dense estimations. The precision of mGHS is the lowest of any model, leading to low accuracy compared to other models. mGHS performs far better in lower dimensional settings (see Supplementary Material 2.5), suggesting poor scalability.
@@
Figure 3: Performance of covdepGE, mGHS, NExON-Bayes & SSJGL in the simulated scenario where there are $P=100$ nodes and $N=200$ samples for each of $\mid\mathcal{A}\mid=4$ networks. $50$ edges ‘appear’ (absolute value of partial correlation linearly increases with covariate) and $50$ edges ‘disappear’ (absolute value of partial correlation linearly decreases with covariate). Performance is assessed over $100$ replicates for mGHS, NExON-Bayes and SSJGL and $10$ replicates for covdepGE due to excessive runtime for large NN (2020+ hours).
Overall, this study exemplifies the advantages of NExON-Bayes in practice. To obtain a more complete assessment of NExON-Bayes’ performance and test its robustness, three further simulation studies are explored that, to a varying degree, deviate from the specification of the model. Whilst the performance of NExON-Bayes worsens in misspecified frameworks (as expected), NExON-Bayes is shown to be robust in these settings by maintaining good performance metrics relative to its competitors. The full analysis of these additional simulation studies can be found in Supplementary Material 2.4.
An analysis of computational runtimes indicates that NExON-Bayes performs inference more quickly than the other joint models that were tested (covdepGE, mGHS, SSJGL) within the range $10\leq P\leq 250$. For up to 8 networks with up to 150 variables ($P\leq 150)$, NExON-Bayes runs in less than 30 minutes and in less than 2 seconds for $P\leq 50$, which is between 1 and 2 orders of magnitude quicker than its competitors. A more complete analysis is given in Supplementary Material 2.5.
@@ 3.4 Real-World Application
We consider proteomic data from The Cancer Genome Atlas (TCGA) (TCGA), specifically, reverse phase protein array data of patients that have been diagnosed with stage I, II and III breast carcinoma [^2] Stage I BRCA refers to tumours that have not spread beyond the breast and are still small in size, stage II refers to slightly larger tumours that may have spread to 1-3 lymph nodes, whilst stage III cancer refers to cancer that has spread to the ‘lymph nodes, the skin of the breast or the chest muscle’ and is larger in size (Mac). (BRCA). These data have been preprocessed using replicate-based normalisation (goldman2020visualizing), a commonly used method for omic data which eliminates unwanted batch effects whilst maintaining approximate Gaussianity, and contain the expression levels of $P=131$ target proteins for $N_{1}=113$ stage I, $N_{2}=429$ stage II and $N_{3}=140$ stage III individuals. We treat tumour stage as our ordinal covariate capturing patient heterogeneity. Although factors like hormonal status also contribute to heterogeneity, we focus on tumour stage as an ordinal covariate to capture broad trends across disease progression. Our aim is to characterize how tumour severity impacts the proteomic network and whether conditional dependencies between protein pairs strengthen or weaken with tumour progression.
We firstly validate the estimations made by NExON-Bayes by consulting the STRING database (szklarczyk2023string), which contains protein-protein interactions that have been experimentally validated to a given confidence level. We calculate the percentage of pathways that are found by NExON-Bayes and SSL that have been validated to a 90% confidence threshold and show that NExON-Bayes finds a higher percentage of these validated pathways than SSL for all three BRCA stages. Another useful analysis is to identify the ‘hub’ proteins, which we define as proteins whose node degree exceeds the 90th percentile, in the networks estimated by NExON-Bayes. This can shed light on which proteins are prominent in which stages of BRCA. For example, NExON-Bayes identifies GATA binding protein 3, Estrogen Receptor $\alpha$ and fibronectin to have degrees in the top 5% in all three stages, whilst Inositol polyphosphate-4-phosphatase (INPP4B) is identified as a hub protein in only the stage III: INPP4B inactivation is known to be common in triple-negative breast cancer cases (Rodgers2021); see Supplementary Material 3.1 & 3.2 for full results of both of these analyses.
@@
NExON-Bayes’ framework provides unique insights into molecular networks by estimating the relation between the covariate and the conditional dependence. In essence, inspecting the edges corresponding to the largest absolute posterior point estimates of $\beta_{ij}$ could help identify biological processes that characterise tumour progression when used to estimate the proteomic networks of BRCA patients. Figure 4 shows the subnetworks consisting of the edges corresponding to the $50$ largest negative values (in absolute terms) and the $50$ largest positive values of the posterior mean estimate of $\beta_{ij}$.
To assess whether the borrowing of information facilitated through the inclusion of the edge-specific $\beta_{ij}$ is beneficial in this setting, we perform a ranked gene set enrichment analysis (GSEA): genes are ranked by the sum of the estimated regression coefficients between the proteins which they encode and all other proteins (with positive and negative networks considered separately as in Figure 2) and the resulting ranked list is then assessed to find whether genes belonging to predefined, biologically curated gene sets are over-represented towards the extremes (top or bottom) of the ranking, indicating non-random enrichment (subramanian2005gene).
We compare the gene sets found to be enriched in the $\beta_{ij}$ coefficients of NExON-Bayes (collectively denoted $\boldsymbol{\beta}_{\text{NExON}}$) with those found to be enriched in the coefficients estimated by GraphR ($\boldsymbol{\beta}_{\text{GraphR}}$), which is a graphical model that accounts for covariate information and adopts a decoupling strategy, wherein $\beta_{ij}$ coefficients are estimated separately before applying a variational Bayes inference algorithm. This differs to NExON-Bayes, which estimates each $\beta_{ij}$ and the precision matrix entries in conjunction within the VBECM algorithm. GraphR also enforces sparsity through a spike-and-slab prior on the precision matrices and uses $\beta_{ij}$ to represent the effect of the covariate on edge-specific regression (chen2025probabilistic). This is very similar to NExON-Bayes and thereby provides a suitable and interesting comparison. GSEA finds 23 gene sets enriched in $\boldsymbol{\beta}_{\text{GraphR}}$, versus 22 in $\boldsymbol{\beta}_{\text{NExON}}$, where a 5% threshold on the nominal $p$-value of each is employed, along with a 25% false discovery rate threshold. (Supplementary Material 3.3, Tables S5 & S6).
The GSEA based on $\boldsymbol{\beta}_{\text{GraphR}}$ identifies several sets related to epidermal growth factor receptor (EGFR), which is a receptor tyrosine kinase often upregulated in various cancers, including breast cancer and colorectal cancer (wee2017epidermal). Also identified are gene sets relating to ErbB protein signaling. ErbB is a member of the EGFR family that is more prominently expressed in metastatic (stage IV) breast cancer than in primary (stage I) tumours (xue2006erbb3). These sets are not found to be enriched in $\boldsymbol{\beta}_{\text{NExON}}$, although colorectal cancer-related pathways are, which could be explained by the confounding link to EGFR regulation.
Several of the gene sets found enriched in $\boldsymbol{\beta}_{\text{NExON}}$ are either directly or indirectly related to apoptosis. The largest of these is the apoptosis modulation and signaling pathway. Apoptosis refers to programmed cell death, a vital process that maintains tissue homeostasis by eliminating damaged or unnecessary cells. In cancer, the evasion of apoptosis allows malignant cells to survive and proliferate uncontrollably. As such, pathways controlling apoptotic modulation and signaling play an important role in cancer progression (plati2011apoptotic).
@@
As a final example, the Ataxia Telangiectasia Mutated (ATM) Protein signaling gene set is found enriched in the positive network of NExON-Bayes. The ATM signaling pathway plays a key role in breast cancer by maintaining genomic stability and influencing treatment responses. As an important regulator of the DNA damage response, ATM mutations can lead to increased susceptibility to malignancy and impact tumor progression (VARADHAN2024155447).
Whilst in this analysis there is no quantitative criterion to ascertain which strategy (coupling vs decoupling) outperforms the other, these results show that both can find several relevant gene sets.
The ability to uncover these gene sets through encapsulating sample-level heterogeneity effectively in a joint model is an important practical advantage of NExON-Bayes (and GraphR). We have shown that, by taking this joint approach, interesting validated or candidate disease mechanisms are uncovered in a disease which progresses through defined stages. These mechanisms warrant dedicated investigation beyond the scope of this paper.
@@ 4 Discussion
We have introduced a new method, NExON-Bayes, that jointly estimates multiple graphical models while accounting for sample-level heterogeneity as encapsulated in auxiliary ordinal covariate information. We show that, for such auxillary data, NExON-Bayes’ formulation better captures sample-level heterogeneity than network estimation models that use continuous covariates or treat pre-defined groups of samples in an exchangeable manner, without resorting to covariate information. Moreover, thanks to its efficient deterministic inference procedure, our approach is more scalable and suited to the analysis of several current real data settings. Applying our novel methodology to the TCGA breast cancer data set reveals new molecular pathways that characterize tumour stage progression, which could lead to new hypotheses about plausible oncogenic mechanisms.
Several avenues for future work could be considered. First, the current prior formulation assumes a linear relationship between the ordinal variable and the probability of edge inclusion – relaxing this linearity assumption through an alternative monotonic parametric formulation may enhance both the model performance and its biological relevance. Secondly, the model could in principle be easily extended to accommodate multiple covariates. We did not pursue this extension here, however, as its practical usefulness is limited: indeed, it would require specifying a separate network for every combination of discrete covariate levels, which would reduce statistical power by fragmenting the sample across many networks. This would also hamper interpretability, as by combining multiple covariates to define networks, they lose their clear ordinality. Next, by using a deterministic inference algorithm, the model does not explicitly represent the full posterior distribution and may therefore underestimate posterior uncertainty (giordano2018covariances). In principle, stochastic inference algorithms can provide more faithful uncertainty quantification, provided convergence is reached. Accordingly, an interesting direction for future work is to investigate inference procedures that yield richer uncertainty quantification, for instance via hybrid deterministic–stochastic approaches, while maintaining a reasonable trade-off between computational cost and inferential accuracy. Another interesting extension concerns the treatment of missing data. Whilst NExON-Bayes does not intrinsically deal with missingness, several imputation methods are available for omics data (lazar2016accounting; qiu2020genomic). A natural extension to this work would be to instead incorporate model-based Bayesian imputation directly within NExON-Bayes by extending the model hierarchy to explicitly describe the observational process. Lastly, as is typical for joint network estimation models, unequal sample sizes can lead to varying levels of sparsity in the estimated networks, reflecting differences in statistical power. This behaviour can in fact be desirable, as smaller sample sizes naturally yield sparser estimates, reducing the risk of false positives. However, if the goal is to isolate true biological variation independent of sample size effects, this pattern can be adjusted for — though doing so will require methodological advances beyond current approaches (das2020nexus).
@@
This research was supported by the Lopez–Loreta Foundation (J.F., H.R., C.L., X.X.) and the UK Medical Research Council programme (J.F. and X.X.). C.L. has also received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No.101126636. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.